I got confused on everything here please help
Below is a preference schedule of a recent election Number of voters 1st choice 2nd choice 3rd choice Answer the 12 O a) What is the total number of votes? O b) Which candidate would win under the plurality method? Oc) What percent of the voters number selected candidate C as their 1t choice? Round to the nearest whole O d) In the one-on-one match up of candidate A versus candidate 8, which candidate is preferred? S. Below is a preference schedule of a recent election t choice 2nd choice 3rd choice Find the points for each candidate using Borda Count. Use the count that assigns 1 point to last place. O a) Candidate A O b) Candidate B O c) Candidate C O d) Candidate D - O e) The is winner is
Answer #1

4.a) The tota number of voters is = (8+7+9+7+11+12) = 54.

b) A has 8+9 = 17 votes.

B has 7+11 = 18 votes.

C has 12 votes.

D has 7 votes.

Therefore, under plurality method, candidate B would win.

c) The required percent of voters selected candidate C as their 1st choice is = (12/54*100)% = 22%.

d) In one-on-one match, A will have 8+9+12 = 29 votes and B will have 7+7+11 = 25 votes.

Therefore, candidate A is preferred.